OpenStudy (anonymous):

how do I convert the numeral to a numeral in base 10? 1110two?

OpenStudy (shamil98):

what do you mean by 1110two?

OpenStudy (anonymous):

the number given is 1110two

OpenStudy (anonymous):

I need to know how to convert it to a numeral in base 10

OpenStudy (anonymous):

how do I do that?

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

Each number in the base 2 number represents that 2 to an increasing power. You multiply that number by the number in that place value. Ex: 1101 = 1(2^3) + 1(2^2) + 0(2^1) +1(2^0) Which is equal to 8 + 4 + 1 which is equal to 13. This works for all numbers.

OpenStudy (anonymous):

so if it asks to convert the numeral to a numeral in base 10 and gives the number 1110two how would I write my answer?

OpenStudy (loser66):

you should go from the right to left, the far right is 0,right? then the first term is 0 * (2^0) = 0 next is 1, right? it is 1* (2^1) = 2 next to the left is 1 again, right? so, it is 1* (2^2) = 4 the last one next to the left of it is 1 again, 1* (2^3) = 8 then add them together ------ = 14 base 10

OpenStudy (loser66):

 pay attention at the exponent, it goes from 0, 1, 2, 3, if you have more , this exponent will move continuously 4,5, 6 and so on